OpenStudy (anonymous):

Solve x + 3 = 11. A. –8 B. 8 C. –14 D. 14

OpenStudy (anonymous):

eliminate x.

OpenStudy (anonymous):

i have 8 question that u need help on will u help me

OpenStudy (shamil98):

x + 3 = 11 A number has 3 added to it, what is that number? So, you would get rid of that number added.

OpenStudy (shamil98):

Correct, it's 8.

OpenStudy (anonymous):

what? he didn't say it?

OpenStudy (shamil98):

oh, i misread that

OpenStudy (anonymous):

Solve 13 – x = 4. A. –17 B. 9 C. –9 D. 17

OpenStudy (shamil98):

x + 3 = 11 3 added to x is 11 so you would subtract the 3 to get x by itself. x + 3 = 11 -3 -3 x = 8

OpenStudy (anonymous):

brb